# UniGraspTransformer: Simplified Policy Distillation for Scalable Dexterous Robotic Grasping

### v1

@@ 1 Introduction
In our experiments, our proposed approach demonstrates substantial improvements over the previous state-of-the-art, UniDexGrasp++ [51], across a range of evaluation settings. Specifically, we evaluate our method in both the state-based setting—where object observations and dexterous hand states are perfectly accurate, as provided by a simulator—and the vision-based setting, where object point clouds are derived from multi-view reconstruction. Our approach consistently outperforms UniDexGrasp++ [51] across multiple object types, including seen objects, unseen objects within seen categories, and entirely unseen objects from unseen categories, as illustrated in Figure 1. For instance, our method achieves performance gains of 3.5%, 7.7%, and 10.1% over UniDexGrasp++ [51] on seen objects, unseen objects within seen categories, and entirely unseen objects from unseen categories, respectively, under the vision-based setting.
@@ 2 Related Works
Dexterous grasping [53, 58, 51, 55, 21, 24, 56, 34, 33] introduces advanced multi-fingered manipulation, enabling more versatile grasps for objects of diverse shapes. However, controlling the highly dexterous multi-fingered system poses significant challenges for traditional analytical techniques [53, 1, 48, 23]. Recent advances have utilized learning-based methods to enable effective dexterous manipulation. One approach decomposes the grasping process into two stages: generating a static grasp pose and then performing a dynamic grasping through trajectory planning or goal-conditioned reinforcement learning [58, 57, 60, 16, 22, 46, 2, 5]. Although promising diversity, the generated static grasp poses are often not validated in dynamic settings, which adversely affects the overall success. Alternatively, another line of approach directly learns the entire grasping process through expert demonstrations from humans or reinforcement learning agents [51, 40, 38, 30, 19, 29, 52]. While effective, these approaches often involve complex training pipelines and experience performance degradation when a single policy is applied across a broad range of objects, due to the limited number of demonstrations and the constrained capacity of student networks.
@@
To overcome these limitations, we extend the latter approach by proposing a novel pipeline that combines online reinforcement learning with large-model offline distillation, simplifying training while improving scalability and grasping performance.
Policy Distillation. Policy distillation [42, 12, 35, 7, 15, 41, 6, 49, 39] provides an effective approach for transferring knowledge from high-performance policies to a single universal policy, promoting both model compactness and generalization across diverse tasks.
In robotics, recent works have focused on combining imitation learning and reinforcement learning [38, 30, 19, 49, 39, 51] to enable student agents to learn from teacher demonstrations. This research generally follows two main approaches based on the source of demonstrations. The first approach directly trains the student policy on pre-collected human demonstrations, such as teleoperated human motions or recorded human videos [38, 30, 19, 10, 4, 43]. While effective, gathering extensive demonstrations can be costly, particularly for complex tasks like dexterous grasping with diverse objects in varied poses, limiting the student’s generalization capacity. The second approach generates demonstrations using pre-trained policies within the generalist-specialist learning framework [15, 11, 31, 47, 58, 51, 52]. Here, the task space is divided into sub-tasks, with reinforcement learning policies specialized and trained for each. These policies are then distilled into a single universal policy, enhancing the agent’s ability to generalize across the full task space. Despite the progress made, a single network handling a broad range of objects often experiences performance drops due to the limited teacher policies and the constrained capacity of student networks, which struggle to capture the complexity of the entire task space.
@@ 3.2 Grasp Trajectory Generation
Each dedicated policy network is now able to grasp its assigned object across various initial poses, achieving an average success rate of 94.1% across the 3,200 training objects. For each object, we randomly rotate it and use its corresponding policy network to generate a successful grasp trajectory. This process is repeated $M$ ($M=1000$ by default) times per object, resulting in a dataset $\mathcal{D}$ of $3,200\times M$ successful grasp trajectories. Each trajectory, $\mathcal{T}=\{(S_{1},A_{1}),\dots,(S_{t},A_{t}),\dots,(S_{T},A_{T})\}$, is a sequence of steps, where $A_{t}$ represents robotic action (18 active DOFs for the fingers and 6 active DOFs for the wrist) at timestep-$t$, and $S_{t}$ represents the information of proprioception, previous action, object state, hand-object distance, time, and object point cloud. Definitions for proprioception, object state, hand-object distance and time have been provided in Table 1. The dataset $\mathcal{D}$ is then used to train our UniGraspTransformer in a supervised manner, as described in Section 3.3.
@@ 3.4 Adaptation to the Vision-Based Setting
Input Adaptation. In the vision-based setting, we use five cameras mounted at the table’s top and borders to estimate the object point clouds. The estimated point clouds consist of two components: 1) the partial object point cloud and 2) the hand point cloud. In our implementation, we segment and remove the hand point cloud, leaving only the partial object point cloud. In contrast, in the state-based setting, the object point clouds are uniformly sampled from the object mesh, which are complete and accurate. This difference affects the input to UniGraspTransformer as follows: 1) for object state representation, we use the center of the partial object point cloud as the object position and apply PCA on this partial cloud to represent the object rotation; 2) we use the partial object point cloud to calculate hand-object distance; and 3) we re-train a dedicated object encoder, termed V-Encoder, to extract features from the partial object points, as will be discussed later. Other configurations, such as network architecture, loss function, and supervision signals, remain unchanged.
@@ 4 Experiments
Implementation Details. For a single dedicated policy network, we create 1,000 simulation environments, and update the policy network every 16 steps over 10K iterations, with a learning rate of 3e-4. Training is conducted on an NVIDIA V100 GPU, requiring approximately 3 hours per policy. The UniGraspTransformer model is trained with a batch size of 800 trajectories over 100 epochs, using a fixed learning rate of 1e-4. Training is performed on 8 NVIDIA A100 GPUs and takes around 70 hours to complete. The object encoders, both state-based and vision-based, are trained on point cloud data with a batch size of 100. Training runs for 800K iterations with a learning rate of 5e-4 on an NVIDIA A100 GPU, requiring about 40 hours.
@@ 4.2 Ablation Studies
Table 5: Analysis of distilling varying numbers of dedicated policy networks (one policy per object) into UniGraspTransformer. The evaluation is conducted on the corresponding object set.
@@
  • Object Number. Table 5 shows the success rate of UniGraspTransformer when distilling varying numbers of object-specific policy networks, evaluated across different object sets with 400, 800, 1,600, and 3,200 objects. The results indicate a high and stable success rate, with a slight decline as the number of objects increases.
@@
Reward Functions for Dedicated Policy Network. As outlined in Section 3.1, the dedicated policies are trained with reward functions defined in Eq.1. We explore two reward functions: (1) the grasp reward $R_{d}$, which penalizes the distance between the dexterous hand and the object, and (2) the reward $R_{o}$, which encourages the hand to stay open until it contacts the object. In our default setup, $R_{d}$ is computed by measuring the distances between 36 hand points and the object point clouds. In Table 9, we examine an alternative version where $R_{d}$ is based on distances between the 36 hand points and the object center. Additionally, Table 9 assesses the impact of including or excluding the reward $R_{o}$. A well-designed reward function enhances the performance of all dedicated policy networks, achieving an average success rate of 94.1% across 3,200 training objects.
@@ B.1 Baseline Methods
Figure 10: Success rates across varying numbers of seen objects.
Figure 11: Comparison of grasp poses. (a) UniDexGrasp++ often grasps different objects using similar poses, characterized by twisted and floated middle fingers. (b) In contrast, UniGraspTransformer adapts to objects of various shapes, exhibiting a diverse array of grasp poses.
@@ Appendix C More Analysis
Qualitative Results. The progressive online distillation approach [15] employed in UniDexGrasp++[51] results in a universal policy that tends to grasp different objects using similar poses, as illustrated in Figure 11. In contrast, our UniGraspTransformer, utilizing a larger model and an offline distillation framework, demonstrates the ability to grasp objects of various shapes with a wide range of diverse poses. This increased diversity in grasping strategies is further highlighted in Figure 12.

### v2

@@ 1 Introduction
In our experiments, our proposed approach demonstrates substantial improvements over the previous state-of-the-art, UniDexGrasp++ [50], across a range of evaluation settings. Specifically, we evaluate our method in both the state-based setting—where object observations and dexterous hand states are perfectly accurate, as provided by a simulator—and the vision-based setting, where object point clouds are derived from multi-view reconstruction. Our approach consistently outperforms UniDexGrasp++ [50] across multiple object types, including seen objects, unseen objects within seen categories, and entirely unseen objects from unseen categories, as illustrated in Figure 1. For instance, our method achieves performance gains of 3.5%, 7.7%, and 10.1% over UniDexGrasp++ [50] on seen objects, unseen objects within seen categories, and entirely unseen objects from unseen categories, under the vision-based setting.
@@ 2 Related Works
Dexterous grasping [52, 57, 50, 54, 19, 22, 55, 34, 33] introduces advanced multi-fingered manipulation, enabling more versatile grasps for objects of diverse shapes. However, controlling the highly dexterous multi-fingered system poses significant challenges for traditional analytical techniques [52, 1, 47, 21]. Recent advances have utilized learning-based methods to enable effective dexterous manipulation. One approach decomposes the grasping process into two stages: generating a static grasp pose and then performing a dynamic grasping through trajectory planning or goal-conditioned reinforcement learning [57, 56, 61, 14, 20, 45, 2, 4]. Although promising diversity, the generated static grasp poses are often not validated in dynamic settings, which adversely affects the overall success. Alternatively, another line of approach directly learns the entire grasping process through expert demonstrations from humans or reinforcement learning agents [50, 40, 38, 29, 17, 28, 51, 60, 59, 23, 30]. While effective, these approaches either require complex training pipelines or suffer significant performance degradation when a single policy is applied across a broad range of objects, due to the limited number of training objects and expert demonstrations, as well as the constrained capacity of their policy networks.
@@
To overcome these limitations, we extend the latter approach by proposing a novel pipeline that integrates online reinforcement learning with large-model offline distillation, simplifying training while improving scalability and grasping performance.
Policy Distillation. Policy distillation [42, 10, 35, 13, 41, 5, 48, 39] provides an effective approach for transferring knowledge from high-performance policies to a single universal policy, promoting both model compactness and generalization across diverse tasks.
In robotics, recent works have focused on combining imitation learning and reinforcement learning [38, 29, 17, 48, 39, 50] to enable student agents to learn from teacher demonstrations. This research generally follows two main approaches based on the source of demonstrations. The first approach directly trains the student policy on pre-collected human demonstrations, such as teleoperated human motions or recorded human videos [38, 29, 17, 8]. While effective, gathering extensive demonstrations can be costly, particularly for complex tasks like dexterous grasping with diverse objects in varied poses, limiting the student’s generalization capacity. The second approach generates demonstrations using pre-trained policies within the generalist-specialist learning framework [13, 9, 31, 46, 57, 50, 51]. Here, the task space is divided into sub-tasks, with reinforcement learning policies specialized and trained for each. These policies are then distilled into a single universal policy, enhancing the agent’s ability to generalize across the full task space. Despite the progress made, a single network handling a broad range of objects often experiences performance drops due to the limited teacher policies and the constrained capacity of student networks, which struggle to capture the complexity of the entire task space.
@@ 3.2 Grasp Trajectory Generation
Each dedicated policy network is now able to grasp its assigned object across various initial poses, achieving an average success rate of 94.1% across the 3,200 training objects. For each object, we randomly rotate it and use its corresponding policy network to generate a successful grasp trajectory. This process is repeated $M$ ($M=1000$) times per object, resulting in a dataset $\mathcal{D}$ of $3,200\times M$ successful grasp trajectories. Each trajectory, $\mathcal{T}=\{(S_{1},A_{1}),\dots,(S_{t},A_{t}),\dots,(S_{T},A_{T})\}$, is a sequence of steps, where $A_{t}$ represents robotic action (18 active DOFs for the fingers and 6 active DOFs for the wrist) at timestep-$t$, and $S_{t}$ represents the observation of proprioception, previous action, object state, hand-object distance, time, and object point cloud, as defined in Table 1. The dataset $\mathcal{D}$ is then used to train our UniGraspTransformer in a supervised manner, as described in Section 3.3.
@@ 3.4 Adaptation to the Vision-Based Setting
Input Adaptation. In the vision-based setting, we use five cameras mounted at the table’s top and borders to estimate the object point clouds. The estimated point clouds consist of two components: 1) the partial object point cloud and 2) the hand point cloud, which is segmented and removed in our implementation. In the state-based setting, the object point clouds are uniformly sampled from the object mesh, which are complete and accurate. This difference affects the input to UniGraspTransformer as follows: 1) for object state representation, we use the center of the partial object point cloud as the object position and apply PCA on this partial cloud to represent the object rotation; 2) we use the partial object point cloud to calculate hand-object distance; and 3) we re-train a dedicated object encoder, termed V-Encoder, to extract features from the partial object points. Other configurations, such as network architecture, loss function, and supervision signals, remain unchanged.
@@ 4 Experiments
Implementation Details. For a single dedicated policy network, we create 1,000 simulation environments, and update the policy network every 16 steps over 10K iterations, with a learning rate of 3e-4. Training is conducted parallelly on 16 NVIDIA V100 GPUs, requiring 80 hours for 3,200 dedicated policies. The UniGraspTransformer model is trained with a batch size of 800 trajectories over 100 epochs, using a fixed learning rate of 1e-4. Training is performed on 8 NVIDIA A100 GPUs and takes around 70 hours to complete. The object encoders, both state-based and vision-based, are trained on point cloud data with a batch size of 100. Training runs for 800K iterations with a learning rate of 5e-4 on an NVIDIA A100 GPU, requiring 40 hours.
@@ 4.2 Ablation Studies
Table 5: Analysis of distilling varying numbers of dedicated policy networks (one policy per object) into UniGraspTransformer. The evaluation is conducted on the corresponding seen object set.
Method DAgger UniGraspTransformer
Success Rate (%) 88.2 92.5
Table 6: Analysis of distilling 400 dedicated policies into one universal policy, via online or offline methods. The evaluation is conducted on the corresponding seen object set.
@@
  • Object Number. Table 5 shows the success rate of our UniGraspTransformer when distilling varying numbers of object-specific policy networks, evaluated across different object sets with 400, 800, 1,600, and 3,200 objects. The results indicate a high and stable success rate, with a slight decline as the number of objects increases.
  • Online vs. Offline. Table 6 compares the online distillation method: DAgger with MLPs, to our offline distillation method: UniGraspTransformer. The results highlight the advantages of offline distillation with large policy networks when handling a diverse set of teacher policies.
@@
Reward Functions for Dedicated Policy Network. As outlined in Section 3.1, the dedicated policies are trained with reward functions defined in Eq.1. We explore two reward functions: (1) the grasp reward $R_{d}$, which penalizes the distance between the dexterous hand and the object, and (2) the reward $R_{o}$, which encourages the hand to stay open until it contacts the object. In our default setup, $R_{d}$ is computed by measuring the distances between 36 hand points and the object point clouds. In Table 10, we examine an alternative version where $R_{d}$ is based on distances between the 36 hand points and the object center. Additionally, Table 10 assesses the impact of including or excluding the reward $R_{o}$. A well-designed reward function enhances the performance of all dedicated policy networks, achieving an average success rate of 94.1% across 3,200 training objects. Please refer to our supplementary material for more details.
@@ B.1 Baseline Methods
Figure 10: Success rates across seen objects.
@@ Appendix C More Analysis
Qualitative Results. The progressive online distillation approach [13] employed in UniDexGrasp++[50] results in a universal policy that tends to grasp different objects using similar poses. In contrast, our UniGraspTransformer, utilizing a larger model and an offline distillation framework, demonstrates the ability to grasp objects of various shapes with a wide range of diverse poses. This increased diversity in grasping strategies is further highlighted in Figure 11.